# VideoNSA: Native Sparse Attention Scales Video Understanding

### v1

@@ Abstract
Video understanding in multimodal language models remains limited by context length: models often miss key transition frames and struggle to maintain coherence across long time scales. To address this, we adapt Native Sparse Attention (NSA) to video-language models. Our method,  VideoNSA, adapts Qwen2.5-VL through end-to-end training on a 216K video instruction dataset. We employ a hardware-aware hybrid approach to attention, preserving dense attention for text, while employing NSA for video. Compared to token-compression and training-free sparse baselines, VideoNSA achieves improved performance on long-video understanding, temporal reasoning, and spatial benchmarks. Further ablation analysis reveals four key findings: (1) reliable scaling to 128K tokens; (2) an optimal global–local attention allocation at a fixed budget; (3) task-dependent branch usage patterns; and (4) the learnable combined sparse attention help induce dynamic attention sinks.
@@ 1 Introduction
Key moments of a video can occur at any time, exemplified by soccer where game deciding moments typically span seconds of a 90 minute game. Within those game deciding moments split second actions define the outcome: an assist, a missed tackle, the movement of the keeper. Multimodal large language models (MLLMs)Team (2025); Team et al. (2025b; a) have achieved substantial progress in vision-language perception and reasoning, but still cannot match humans ability to extract and reason about salient moments in videos. While humans naturally sample color visuals around 60hz, (Kalloniatis & Luu, 2007) across large contexts, existing VLMs often sample a single frame per second. Intuitively, increasing the context for these models by sampling more frames improves accuracy Cai et al. (2024); Wu et al. (2024), particularly for long videos and complex reasoning tasks. However, this approach pays for improvement with additional tokens, increasing computational complexity and pushing against fundamental limits of model context.
To address these challenges, many approaches Wang et al. (2024); Li et al. (2024b); Jin et al. (2024); Wang et al. (2025a); Yang et al. (2024) adopt token compression to reduce redundancy and increase informative context. However, when applied to complex reasoning tasks, these compression-based models perform worse compared to full-token methods Song et al. (2025a). Moreover, compression strategies often limit generalization through reduced perception and reasoning capacity Wen et al. (2025). In contrast, sparse attention mechanisms preserve tokens, but focus the models capabilities on relevant dependencies between tokens. Numerous sparse attention methods have already been employed in large language models (LLMs), but most are inadequate for video complexity (detailed in Appendix A). Therefore, we present VideoNSA, which adopts Native Sparse Attention Yuan et al. (2025b), a learnable hardware-aware sparse attention mechanism proven to be effective in long-context modeling. VideoNSA is the first learnable and hardware-aware sparse attention framework tailored for video understanding, effectively scaling to ultra-long vision-text context. We apply the learnable sparse attention to video token sequences, while preserving grouped-query attention for text tokens. Following this pattern, our experiments show that using only 3.6% of the attention budget on 128K context length while improving performance on various tasks
@@ 2.1 Preliminaries
Existing training-free sparse attention methods are rarely hardware aligned, and typically don’t increase training efficiency. Native Sparse Attention Yuan et al. (2025b) (NSA) avoids computing attention between all key-value pairs $(\mathbf{K}_{t},\mathbf{V}_{t})$, instead, for each query $\mathbf{q}_{t}$, NSA dynamically constructs an information-dense KV cache subset. NSA combines three complementary cache branches with a learnable gate $g_{t}^{c}$ adaptively weighting each branch yielding $\mathbf{o}_{t}$:
@@ 2.2 Architecture
Existing token compression methods Yang et al. (2025d); Zhang et al. (2025c); Hyun et al. (2025); Zhang et al. (2025h) suffer from irreversible information loss on complex tasks and don’t address computational and latency bottlenecks in LLM video understanding. From the perspective of attention as a message passing in a Graph Neural Network Joshi (2025); Pappone (2025), it’s clear this bottleneck is fundamental. Standard attention propagates information between nodes (tokens) through edges (attention weights), with each token being updated by aggregating features from its neighbors, weighted by attention scores. Training-free sparse attention often imposes a static adjacency matrix whose fixed subgraph connectivity restricts information flow. Conversely, NSA Yuan et al. (2025b) provides data-dependent sparsity that preserves edges necessary for a particular task.
We build VideoNSA upon Qwen2.5-VL-7B Qwen et al. (2025), which incorporates a vision encoder and adopts Qwen2.5-7B Bai et al. (2025) as the LLM. As illustrated in Figure 1, VideoNSA introduces a hybrid attention mechanism in the LLM across different modalities. At each layer $l$, we split the input tokens $\mathbf{X}^{(l-1)}$ into vision tokens $\mathbf{X}^{(l-1)}_{\mathcal{V}}$ and text tokens $\mathbf{X}^{(l-1)}_{\mathcal{T}}$ according to their position IDs. For vision tokens, VideoNSA applies NSA Yuan et al. (2025b) with a dedicated gate $g^{c}_{t}$ on each head. We set the block size $s$ equal to the token number per frame, and obtain the block-level representation by averaging all tokens within the block. The vision attention output $o_{\mathcal{V}}$ is dynamically weighted by the compression, selection, and sliding window branches as:
@@ 2.3 Training Recipe
We conduct end-to-end training to adapt vision features for data-dependent sparse connectivity in the language model. The training dataset of VideoNSA is constructed from LLaVA-Video-178K Zhang et al. (2024c) by filtering for question answer pairs at 4 fps and retaining videos with 350–550 frames, for a subset of 216K pairs. To emphasize sparse attention for temporal redundancy, we constrain the maximum pixels per frame to 50,176, and the maximum context length per training instance to 36K tokens. In VideoNSA, block size $s$ is set to 64, block $b$ is set to 32, and sliding window size $w$ is set to 256. We trained using SWIFT Zhao et al. (2024), adapting the NSA Yuan et al. (2025b) implementation from FLA Yang & Zhang (2024) and Pai et al. (2025b). The complete training process requires 4600 H100 GPU hours. More training details including hyper-parameters selection can be found in Appendix B.
@@ 3.1 Effectiveness on Video Understanding
Table 1: Results on long video understanding, temporal reasoning and spatial understanding tasks. LVB, LTS for LongVideobench Wu et al. (2024) and LongTimeScope Zohar et al. (2025).
@@
LLaVA-Video-7B Zhang et al. (2024b) 58.2 74.1 34.0 35.6
@@
We evaluate VideoNSA across three domains including long video understanding, temporal reasoning, and spatial understanding using LMMs-Eval Zhang et al. (2024a) and VLMEvalKit Duan et al. (2024). Table 1 indicates that sparse attention methods consistently outperform token compression approaches. We empirically evaluate the effectiveness of VideoNSA based on several popular long video understanding benchmarks, including LongVideoBench Wu et al. (2024), MLVU Zhou et al. (2024), TimeScope Zohar et al. (2025) and LongTimeScope Zohar et al. (2025). VideoNSA achieves competitive results, narrowing the gap with state-of-the-art methods. We observe that VideoNSA shows clear advantages on tasks involving order-sensitive temporal reasoning and ultra-long video settings (10 hours in LongTimeScope Zohar et al. (2025)). To evaluate the visual temporal reasoning capbility of VideoNSA, we evaluate VideoNSA on Tomato Shangguan et al. (2024), a benchmark spanning six reasoning types and three video scenarios. VideoNSA attains the highest accuracy on Tomato Shangguan et al. (2024), substantially outperforming compression-based methods, underscoring their limitations in fine-grained temporal inference. VSIBench Yang et al. (2025a) focuses on spatial reasoning allowing us to test whether efficient models can preserve local fidelity while achieving efficiency.  VideoNSA matches the strongest sparse attention baselines and significantly surpasses token compression methods in spatial understanding, confirming that it preserves spatial fidelity. All detailed evaluation settings and subset results can be found in Appendix C, Appendix D, Appendix E, and Appendix F.
@@ 3.2 Ablation Study
To further analyze the components of VideoNSA, we visualize attention pattern in each branch in Appendix G and assess the effectiveness of different branches. Table 2 shows that single-branch models suffer significant degradation, and even two-branch combinations remain inferior to the full VideoNSA, highlighting the necessity of integrating all three branches with dynamic gating. Detailed results of different branch combination can be found in Appendix H.
Table 2: Ablation study on branch selection across different tasks. LVB, LTS for LongVideobench Wu et al. (2024) and LongTimeScope Zohar et al. (2025).
@@ 4 Scaling Analysis and Findings
We further examine whether the learned QKV weights of VideoNSA  can imrpove performance in dense attention inference. Table 3 reports the relative performance change over the Qwen2.5-VL-7B Qwen et al. (2025). Due to the limited quality of the training data, our fine-tuned Qwen2.5-VL-7B (Dense-SFT) exhibits slight performance drops on most benchmarks. We observe that the transferred model (Dense-NSA) allows the dense variant to recover and surpass the baseline on several benchmarks suggesting that sparse-trained weights provides inductive bias towards more effective attention distributions. However, the effect remains limited on LongVideoBench Wu et al. (2024). VideoNSA significantly outperforms Dense-NSA on most tasks, highlighting the importance of runtime sparsity and dynamic gating.
@@
The effective vision context length $L$ is jointly determined by the number of vision tokens per frame $T$ and the total number of input frames $F$. VideoNSA is trained with a maximum context length of $L=36K$ tokens, corresponding to $T=64$ tokens per frame. We conduct an information budget study under a fixed context length, by varying tokens per frame and frame rate. We then scale up the context length beyond the training budget, evaluating up to the maximum 128K tokens supported by the language model. As observed in Figure 2, the model consistently achieves higher performance when scaled to longer contexts beyond its training length across benchmarks. However, the ideal allocation of same token budget is highly task-dependent. LongVideoBench Wu et al. (2024) favors allocating more tokens per frame, while Tomato Shangguan et al. (2024) and TimeScope Zohar et al. (2025) benefit more from increasing the number of frames, emphasizing temporal coverage. VSIBench Yang et al. (2025a) shows mixed preferences depending on context length, reflecting a balance between spatial and temporal sampling. Additional results on information scaling are reported in Appendix I.
@@
To determine the optimal attention allocation, we first fix the total sequence length $L$, the attention budget $K_{\text{vis}}$, and the block size $S=64$, while systematically varying the local attention ratio $\alpha=\frac{w}{K_{\text{attn}}}$. We then employ the optimal allocation ratio $\alpha^{\star}$ for attention budget scaling. As shown in Figure 3, scatter points denote different allocation strategies, with their size and color reflecting performance. We highlight the point corresponding to the training configuration, connect equal-budget settings with solid orange lines, and extend the best-performing configuration with dashed lines, where the annotated values indicate the fraction of attention used $\gamma$. Results show that model performance is highly sensitive to attention allocation. Although the optimal ratio between global and local attention varies across tasks, configurations close to the training allocation generally yield better results. Under the same budget, fine-tuning around the training setting often improves performance, whereas simply enlarging the overall budget does not consistently bring further gains. Moreover, across most benchmarks, increasing global attention (enlarging the block count) tends to outperform increasing local attention (enlarging the sliding window). Remarkably, VideoNSA achieves leading performance using only $3.6\%$ of the full attention budget. More results are in Appendix J.
@@
We analyze the gating distribution of VideoNSA across Tomato Shangguan et al. (2024), VSIBench Yang et al. (2025a), and LongVideoBench Wu et al. (2024), and aggregate the average routing gate weights over 100 examples from each. As illustrated in Figure 5, where shaded bars denote the interquartile range and horizontal lines represent mean values, each head in VideoNSA exhibits distinct and diverse preferences across branches throughout its full depth. The diversity allows different layers to specialize in distinct modes of the context-dependent information flow. The compression branch maintains relatively high average weights across most layers, underscoring its primary role in reducing redundancy while preserving salient features. The selection and sliding window gates fluctuate more strongly, occasionally surpassing the compression branch in early and middle layers. However, their contributions diminish in the final layers (e.g., L22–L26), demonstrating that the focus shifts towards aggregating high-level features. We also note strange behavior in the last layer, where all three branches are fully active despite selection and sliding window being inactive in the layers before. Full gate values distribution in Appendix K.
We further dive into the inter-head gate similarity of each layer in Figure 5. In the middle layers, both selection and sliding window gates exhibit pronounced increases in inter-head similarity. This indicates that multiple mid-layer heads converge to highly consistent gating behaviors when the model performs block selection and local temporal integration. However, the compression gate shows consistently low inter-head similarity, indicating that it operates largely in a head-independent manner. At both the initial and final layers of VideoNSA, inter-head similarity remains weak across all gates, reflecting the need to maintain diversity in early representations and to support mixing information in higher-level abstractions. More inter-head gate similarites visualization in Appendix L.
@@
Figure 9 indicates that VideoNSA maintains low sink ratios overall, with only minor fluctuations across layers. However, Flash Attention exhibits a gradual increase in sink ratios toward deeper layers. The compression branch maintains relatively high sink levels across most layers. The selection branch remains consistently close to zero, while the sliding window branch occasionally shows higher peaks in the middle-to-late layers, indicating that locality constraints may still introduce bias in long-sequence settings. From the perspective of positional distribution in Figure 9, Flash Attention produces sinks that are uniformly spread across the entire sequence due to its fully connected dense attention. Under dynamic gating,VideoNSA achieves smoother temporal coverage, alleviating over-reliance on early positions while avoiding the global diffusion characteristic of dense attention. In contrast, the compression branch exhibits strong accumulation at the beginning with an even steeper decay, indicating that token merging exerts its strongest impact on early-stage representations. The selection branch yields very few sinks across the sequence, while the sliding window branch produces sparse peaks at periodic boundaries of local neighborhoods. More analysis about attention sinks on various sparse attention settings can be found in Appendix M.

### v2

@@ Abstract
Video understanding in multimodal language models remains limited by context length: models often miss key transition frames and struggle to maintain coherence across long time scales. To address this, we adapt Native Sparse Attention (NSA) to video-language models. Our method,  VideoNSA, adapts Qwen2.5-VL through end-to-end training on a 216K video instruction dataset. We employ a hardware-aware hybrid approach to attention, preserving dense attention for text, while employing NSA for video. Compared to token-compression and training-free sparse baselines, VideoNSA achieves competitive performance on long-video understanding, temporal reasoning, and spatial benchmarks. Further ablation analysis reveals four key findings: (1) reliable scaling to 128K tokens; (2) an optimal global–local attention allocation at a fixed budget; (3) task-dependent branch usage patterns; and (4) the learnable combined sparse attention help induce dynamic attention sinks.
@@ 1 Introduction
Key moments of a video can occur at any time, exemplified by soccer where game deciding moments typically span seconds of a 90 minute game. Within those game deciding moments split second actions define the outcome: an assist, a missed tackle, the movement of the keeper. Multimodal large language models (MLLMs)Team (2025); Team et al. (2025b; a); Qiu et al. (2025); Yuan et al. (2025a); Gui et al. (2025); Chung et al. (2025); Li et al. (2025b) have achieved substantial progress in vision-language perception and reasoning, but still cannot match humans ability to extract and reason about salient moments in videos. While humans naturally sample color visuals around 60hz, (Kalloniatis & Luu, 2007) across large contexts, existing VLMs often sample a single frame per second. Intuitively, increasing the context for these models by sampling more frames improves accuracy Cai et al. (2024); Wu et al. (2024a), particularly for long videos and complex reasoning tasks. However, this approach pays for improvement with additional tokens, increasing computational complexity and pushing against fundamental limits of model context.
To address these challenges, many approaches Wang et al. (2024); Li et al. (2024b); Jin et al. (2024); Wang et al. (2025a); Yang et al. (2024) adopt token compression to reduce redundancy and increase informative context. However, when applied to complex reasoning tasks, these compression-based models perform worse compared to full-token methods Song et al. (2025a). Moreover, compression strategies often limit generalization through reduced perception and reasoning capacity Wen et al. (2025). In contrast, sparse attention mechanisms preserve tokens, but focus the models capabilities on relevant dependencies between tokens. Numerous sparse attention methods have already been employed in large language models (LLMs), but most are inadequate for video complexity (detailed in Appendix A). Therefore, we present VideoNSA, which adopts Native Sparse Attention Yuan et al. (2025c), a learnable hardware-aware sparse attention mechanism proven to be effective in long-context modeling. VideoNSA is the first learnable and hardware-aware sparse attention framework tailored for video understanding, effectively scaling to ultra-long vision-text context. We apply the learnable sparse attention to video token sequences, while preserving grouped-query attention for text tokens. Following this pattern, our experiments show that using only 3.6% of the attention budget on 128K context length while improving performance on various tasks
@@ 2.1 Preliminaries
Existing training-free sparse attention methods are rarely hardware aligned, and typically don’t increase training efficiency. Native Sparse Attention Yuan et al. (2025c) (NSA) avoids computing attention between all key-value pairs $(\mathbf{K}_{t},\mathbf{V}_{t})$, instead, for each query $\mathbf{q}_{t}$, NSA dynamically constructs an information-dense KV cache subset. NSA combines three complementary cache branches with a learnable gate $g_{t}^{c}$ adaptively weighting each branch yielding $\mathbf{o}_{t}$:
@@ 2.2 Architecture
Existing token compression methods Yang et al. (2025d); Zhang et al. (2025c); Hyun et al. (2025); Zhang et al. (2025h) suffer from irreversible information loss on complex tasks and don’t address computational and latency bottlenecks in LLM video understanding. From the perspective of attention as a message passing in a Graph Neural Network Joshi (2025); Pappone (2025), it’s clear this bottleneck is fundamental. Standard attention propagates information between nodes (tokens) through edges (attention weights), with each token being updated by aggregating features from its neighbors, weighted by attention scores. Training-free sparse attention often imposes a static adjacency matrix whose fixed subgraph connectivity restricts information flow. Conversely, NSA Yuan et al. (2025c) provides data-dependent sparsity that preserves edges necessary for a particular task.
We build VideoNSA upon Qwen2.5-VL-7B Qwen et al. (2025), which incorporates a vision encoder and adopts Qwen2.5-7B Bai et al. (2025) as the LLM. As illustrated in Figure 1VideoNSA introduces a hybrid attention mechanism in the LLM across different modalities. At each layer $l$, we split the input tokens $\mathbf{X}^{(l-1)}$ into vision tokens $\mathbf{X}^{(l-1)}_{\mathcal{V}}$ and text tokens $\mathbf{X}^{(l-1)}_{\mathcal{T}}$ according to their position IDs. For vision tokens, VideoNSA applies NSA Yuan et al. (2025c) with a dedicated gate $g^{c}_{t}$ on each head. We set the block size $s$ equal to the token number per frame, and obtain the block-level representation by averaging all tokens within the block. The vision attention output $o_{\mathcal{V}}$ is dynamically weighted by the compression, selection, and sliding window branches as:
@@ 2.3 Training Recipe
We conduct end-to-end training to adapt vision features for data-dependent sparse connectivity in the language model. The training dataset of VideoNSA is constructed from LLaVA-Video-178K Zhang et al. (2024d) by filtering for question answer pairs at 4 fps and retaining videos with 350–550 frames, for a subset of 216K pairs. To emphasize sparse attention for temporal redundancy, we constrain the maximum pixels per frame to 50,176, and the maximum context length per training instance to 36K tokens. In VideoNSA, block size $s$ is set to 64, block $b$ is set to 32, and sliding window size $w$ is set to 256. We trained using SWIFT Zhao et al. (2024), adapting the NSA Yuan et al. (2025c) implementation from FLA Yang & Zhang (2024) and Pai et al. (2025b). The complete training process requires 4600 H100 GPU hours. More training details including hyper-parameters selection can be found in Appendix B.
@@ 3.1 Effectiveness on Video Understanding
Table 1: Results on long video understanding, temporal reasoning and spatial understanding tasks. LVB, LTS for LongVideobench Wu et al. (2024a) and LongTimeScope Zohar et al. (2025).
@@
LLaVA-Video-7B Zhang et al. (2024c) 58.2 74.1 34.0 35.6
@@
We evaluate VideoNSA across three domains including long video understanding, temporal reasoning, and spatial understanding using LMMs-Eval Zhang et al. (2024a) and VLMEvalKit Duan et al. (2024). Table 1 indicates that sparse attention methods consistently outperform token compression approaches. We empirically evaluate the effectiveness of VideoNSA based on several popular long video understanding benchmarks, including LongVideoBench Wu et al. (2024a), MLVU Zhou et al. (2024), TimeScope Zohar et al. (2025) and LongTimeScope Zohar et al. (2025)VideoNSA achieves competitive results, narrowing the gap with state-of-the-art methods. We observe that VideoNSA shows clear advantages on tasks involving order-sensitive temporal reasoning and ultra-long video settings (10 hours in LongTimeScope Zohar et al. (2025)). To evaluate the visual temporal reasoning capbility of VideoNSA, we evaluate VideoNSA on Tomato Shangguan et al. (2024), a benchmark spanning six reasoning types and three video scenarios. VideoNSA attains the highest accuracy on Tomato Shangguan et al. (2024), substantially outperforming compression-based methods, underscoring their limitations in fine-grained temporal inference. VSIBench Yang et al. (2025a) focuses on spatial reasoning allowing us to test whether efficient models can preserve local fidelity while achieving efficiency.  VideoNSA matches the strongest sparse attention baselines and significantly surpasses token compression methods in spatial understanding, confirming that it preserves spatial fidelity. All detailed evaluation settings and subset results can be found in Appendix C, Appendix D, Appendix E, and Appendix F.
@@ 3.2 Ablation Study
To further analyze the components of VideoNSA, we visualize attention pattern in each branch in Appendix H and assess the effectiveness of different branches. Table 2 shows that single-branch models suffer significant degradation, and even two-branch combinations remain inferior to the full VideoNSA, highlighting the necessity of integrating all three branches with dynamic gating. Detailed results of different branch combination can be found in Appendix I.
Table 2: Ablation study on branch selection across different tasks. LVB, LTS for LongVideobench Wu et al. (2024a) and LongTimeScope Zohar et al. (2025).
@@ 4 Scaling Analysis and Findings
We further examine whether the learned QKV weights of VideoNSA  can imrpove performance in dense attention inference. Table 3 reports the relative performance change over the Qwen2.5-VL-7B Qwen et al. (2025). Due to the limited quality of the training data, our fine-tuned Qwen2.5-VL-7B (Dense-SFT) exhibits slight performance drops on most benchmarks. We observe that the transferred model (Dense-NSA) allows the dense variant to recover and surpass the baseline on several benchmarks suggesting that sparse-trained weights provides inductive bias towards more effective attention distributions. However, the effect remains limited on LongVideoBench Wu et al. (2024a). VideoNSA significantly outperforms Dense-NSA on most tasks, highlighting the importance of runtime sparsity and dynamic gating.
@@
The effective vision context length $L$ is jointly determined by the number of vision tokens per frame $T$ and the total number of input frames FFVideoNSA is trained with a maximum context length of $L=36K$ tokens, corresponding to $T=64$ tokens per frame. We conduct an information budget study under a fixed context length, by varying tokens per frame and frame rate. We then scale up the context length beyond the training budget, evaluating up to the maximum 128K tokens supported by the language model. As observed in Figure 2, the model consistently achieves higher performance when scaled to longer contexts beyond its training length across benchmarks. However, the ideal allocation of same token budget is highly task-dependent. LongVideoBench Wu et al. (2024a) favors allocating more tokens per frame, while Tomato Shangguan et al. (2024) and TimeScope Zohar et al. (2025) benefit more from increasing the number of frames, emphasizing temporal coverage. VSIBench Yang et al. (2025a) shows mixed preferences depending on context length, reflecting a balance between spatial and temporal sampling. Additional results on information scaling are reported in Appendix J.
@@
To determine the optimal attention allocation, we first fix the total sequence length $L$, the attention budget $K_{\text{vis}}$, and the block size $S=64$, while systematically varying the local attention ratio $\alpha=\frac{w}{K_{\text{attn}}}$. We then employ the optimal allocation ratio $\alpha^{\star}$ for attention budget scaling. As shown in Figure 3, scatter points denote different allocation strategies, with their size and color reflecting performance. We highlight the point corresponding to the training configuration, connect equal-budget settings with solid orange lines, and extend the best-performing configuration with dashed lines, where the annotated values indicate the fraction of attention used $\gamma$. Results show that model performance is highly sensitive to attention allocation. Although the optimal ratio between global and local attention varies across tasks, configurations close to the training allocation generally yield better results. Under the same budget, fine-tuning around the training setting often improves performance, whereas simply enlarging the overall budget does not consistently bring further gains. Moreover, across most benchmarks, increasing global attention (enlarging the block count) tends to outperform increasing local attention (enlarging the sliding window). Remarkably, VideoNSA achieves leading performance using only $3.6\%$ of the full attention budget. More results are in Appendix L.
@@
We analyze the gating distribution of VideoNSA across Tomato Shangguan et al. (2024), VSIBench Yang et al. (2025a), and LongVideoBench Wu et al. (2024a), and aggregate the average routing gate weights over 100 examples from each. As illustrated in Figure 5, where shaded bars denote the interquartile range and horizontal lines represent mean values, each head in VideoNSA exhibits distinct and diverse preferences across branches throughout its full depth. The diversity allows different layers to specialize in distinct modes of the context-dependent information flow. The compression branch maintains relatively high average weights across most layers, underscoring its primary role in reducing redundancy while preserving salient features. The selection and sliding window gates fluctuate more strongly, occasionally surpassing the compression branch in early and middle layers. However, their contributions diminish in the final layers (e.g., L22–L26), demonstrating that the focus shifts towards aggregating high-level features. We also note strange behavior in the last layer, where all three branches are fully active despite selection and sliding window being inactive in the layers before. Full gate values distribution in Appendix N.
We further dive into the inter-head gate similarity of each layer in Figure 5. In the middle layers, both selection and sliding window gates exhibit pronounced increases in inter-head similarity. This indicates that multiple mid-layer heads converge to highly consistent gating behaviors when the model performs block selection and local temporal integration. However, the compression gate shows consistently low inter-head similarity, indicating that it operates largely in a head-independent manner. At both the initial and final layers of VideoNSA, inter-head similarity remains weak across all gates, reflecting the need to maintain diversity in early representations and to support mixing information in higher-level abstractions. More inter-head gate similarites visualization in Appendix O.
@@
Figure 9 indicates that VideoNSA maintains low sink ratios overall, with only minor fluctuations across layers. However, Flash Attention exhibits a gradual increase in sink ratios toward deeper layers. The compression branch maintains relatively high sink levels across most layers. The selection branch remains consistently close to zero, while the sliding window branch occasionally shows higher peaks in the middle-to-late layers, indicating that locality constraints may still introduce bias in long-sequence settings. From the perspective of positional distribution in Figure 9, Flash Attention produces sinks that are uniformly spread across the entire sequence due to its fully connected dense attention. Under dynamic gating,VideoNSA achieves smoother temporal coverage, alleviating over-reliance on early positions while avoiding the global diffusion characteristic of dense attention. In contrast, the compression branch exhibits strong accumulation at the beginning with an even steeper decay, indicating that token merging exerts its strongest impact on early-stage representations. The selection branch yields very few sinks across the sequence, while the sliding window branch produces sparse peaks at periodic boundaries of local neighborhoods. More analysis about attention sinks on various sparse attention settings can be found in Appendix S.